OpenStudy (anonymous):

solve for y -4(24+8y)=-64.... if you give me thwe anwser I'll give you a metal

OpenStudy (anonymous):

-96-32y=-64 -32y=32 y=-1

OpenStudy (anonymous):

-96- 32y=-64 -96+64= 32y -32= 32y y= -1